# Alphazero-like Tree-Search can guide large language model decoding and training

### v1

@@ 3.2.2 Tree Search Algorithms
Given a learned value function, in this section, we present five types of tree-search algorithms. Due to the space limitation, we leave the background and preliminaries in Appendix C.
MCTS with value function approximation (named as MCTS-$\alpha$): This is the MCTS variants utilized in AlphaZero (Silver et al., 2017a). Starting from the initial state, we choose the node of state $s_{t}$ as the root node and do $n$ times of search simulations consisting of select, expand and evaluate and backup, where the leaf node value evaluated by the learned value function will be backpropagated to all its ancestor nodes. After the search, we choose an action proportional to the root node’s exponentiated visit count, i.e. $a\sim\frac{N(s_{t},a)^{1/\tau}}{\sum_{b}N(s_{t},b)^{1/\tau}}$, and move to the corresponding next state. The above iteration will be repeated until finished. MCTS-$\alpha$ has two main features. Firstly, MCTS-$\alpha$ cannot trace back to its previous states once it takes an action. So it cannot restart the search from the initial state unless multiple searches are conducted which will be discussed in Section 3.2.3. Secondly, MCTS-$\alpha$ utilizes a value function so it can conduct the backward operation during the intermediate steps, without the need to complete the full generation to obtain a Monte-Carlo estimate.
MCTS: This approach was adopted in RAP (Hao et al., 2023), which refers to classic MCTS(Kocsis & Szepesvári, 2006). Specifically, the process starts from the initial state node and then select until finding a leaf node, if this node is non-terminal, expand it and repeat the previous select and expand steps. If the process encounters a terminal node, its value will be evaluated and backup to the whole path. In contrast to MCTS-$\alpha$, it only back-propagates the value on the terminal nodes, relying on a Monte-Carlo estimate of value, and it always starts searching from the initial state node.
MCTS-Rollout: Combining the features from the two algorithms above, we propose a new variant MCTS-Rollout for tree search. Similar to MCTS, MCTS-Rollout always starts from the initial state node. It further does the search simulations analogous to MCTS-$\alpha$, and the backup process can happen in the intermediate step with value function. It repeats the operations above until the process finds $k$ complete answers or reaches the computation limit (e.g. maximum number of tokens.)
@@
Breadth-first/Depth-first Search (BFS/DFS): These two search algorithms were adopted in ToT (Yao et al., 2023). The core idea is to utilize the value function to prune the tree for efficient search, while such pruning happens in tree breadth and tree depth respectively. BFS can be regarded as a beam-search with cumulative reward as the objective. In BFS, nodes with $k$-largest values are maintained and expanded in each step. DFS chooses an unvisited node with the highest value until it finds a complete path. To improve the search efficiency, DFS prunes trees by either dropping the nodes whose value is below a predefined threshold or dropping the nodes of the $n$-lowest value.
@@ 3.2.3 Multiple Search and Search Aggregation
Inspired by Wang et al. (2022) and Uesato et al. (2022) that LLM can improve its performance on reasoning tasks by sampling multiple times and aggregating the candidates, TS-LLM also have the potential to aggregate answers generated by multiple tree searches. There are mainly two choices:
Intra-tree Search. Intra-tree search conducts multiple tree searches on exactly the same tree. Such a method is computationally efficient as the search tree can be reused multiple times. However, the diversity of multiple generations might decrease because the former tree search might influence the latter tree searches. Also, the search space is limited by the sentence-level actions which are previously expanded in the tree and are fixed across multiple tree searches.
Inter-tree Search. Inter-tree Search builds a new tree for each new search, which increases the diversity of the search space, with extra computation burdens proportional to the search times.
Our next step is to aggregate these search results to obtain the final answer. With a learned ORM, we mainly consider the following three different aggregation methods:
Majority Vote. Wang et al. (2022) aggregates the answer $f$ by using majority vote over multiple random generation $y_{i}$: $f^{*}=\arg\max_{f}\sum\nolimits_{\mathbf{y}^{k}}\mathbf{1}_{\text{final\_ans}(\mathbf{y}^{k})=f}$, where $\mathbf{1}$ is the indicator function.
ORM-Max. Given an outcome reward model, the aggregation can choose the answer $f$ with maximum final reward, $f^{*}=\text{final\_ans}(\arg\max_{\mathbf{y}^{k}}\hat{r}_{\theta}(\mathbf{y}^{k}|\mathbf{x}^{k}))$.
ORM-Vote. Given an outcome reward model, the aggregation can choose the answer $f$ with the sum of reward, namely $f^{*}=\arg\max_{f}\sum\nolimits_{\mathbf{y}^{k};\text{final\_ans}(\mathbf{y}^{k})=f}\hat{r}_{\theta}(\mathbf{y}^{k}|\mathbf{x}^{k})$
@@ 3.3 Enhancing LLM Training with Tree Search
By plugging in the tree search operation, LLM can decode better compared to its original decoding performance during inference, which in other words, is a policy improvement process. Based on this, a new training/finetuning paradigm can be proposed, consisting of the following iterative process:
Policy improvement: We firstly generate tree-search trajectories based on $\pi_{\theta_{\text{old}}}$, $v_{\theta_{\text{old}}}$, and $\hat{r}_{\theta_{\text{old}}}$. By imitating new trajectories, LLM policy can be further improved to $\pi_{\theta_{\text{new}}}$.
Policy evaluation: Train value function $v_{\theta_{\text{new}}}$ and ORM $\hat{r}_{\theta_{\text{new}}}$ over the new sampled trajectories.
Such iterative process belongs to generalized policy iteration (Sutton & Barto, 2018), which is also the procedure used in AlphaZero’s training. In our case, the training process involves finetuning three networks on the tree-search augmented training dataset.: (1) Policy network $\pi_{\theta}$: Use cross-entropy loss with trajectories’ tokens as target (2) Value network $v_{\theta}$: Mean squared error loss with trajectories’ temporal difference (TD) or Monte-Carlo (MC) based value estimation as target, and (3) ORM $\hat{r_{\theta}}$: Mean squared error loss with trajectories’ final reward as target.
@@ 3.4 Tree Search’s extra Computation Burdens
Tree-search algorithms will inevitably bring in additional computation burdens, especially in the node expansion phase for calculating legal child nodes and their corresponding value. Prior methodologies, such as ToT and RAP, tend to benchmark their performance against baseline algorithms using an equivalent number of generation paths (named Path@N). This approach overlooks the additional computational demands of the tree-search process and makes the comparison unfair.
In the setting of sentence-level action nodes, such computational demands are substantial. This is primarily because expanding a tree requires the generation of multiple sentences from a single parent node. In contrast, a token-level action node can alleviate the computation demand for all child nodes. A single forward inference can directly yield a probability distribution for all tokens so its search space and computational overhead align closely with the random sampling strategy typically used in LLM generation. Despite this, additional exploration steps (especially in Alphazero-like MCTS search) still introduce further computation. To enable a more fair comparison between tree-search and other baselines, we also should monitor the number of tokens generated for node expansion in our experiments. This helps us juxtapose the performance of different algorithms when operating under comparable token generation conditions.
@@ 4 Experiments
In this section, we first introduce detailed experimental setups in Section 4.1. In Section 4.2, we comprehensively present features of TS-LLM by answering six questions. Refer to Appendix D and E for more settings, implementations, results, and visualizations. Our code is open-sourced at https://github.com/waterhorse1/LLM_Tree_Search.
@@ 4.1 Experiment Setups
Table 1: Task setups
Task Category Tree Max width Tree Max depth Node Train/test size
GSM8k Mathematical Reasoning 6 8 Sentence 7.5k / 1.3k
Game24 Mathematical Planning 20 4 Sentence 1.0k / 0.3k
PrOntoQA Logical Reasoning 6 15 Sentence 4.5k / 0.5k
RLHF Alignment 50 64 Token 30k / 3k
Task setups. For a given MDP, the nature of the search space is primarily characterized by two dimensions: depth and width. To showcase the efficacy of tree-search algorithms across varied search spaces, we evaluate all algorithms on four tasks with different search widths and depths, including the mathematical reasoning task GSM8k (Cobbe et al., 2021), mathematical planning task Game24 (Yao et al., 2023), logical reasoning task PrOntoQA (Saparov & He, 2022), and RLHF alignment task using synthetic RLHF data (Dahoas, ). The specific task statistics and search space hyperparameters are listed in Table 1. These hyperparameters, especially max search width and search depth, are determined by our exploration experiments. They can effectively present the characteristics of a task’s search space. Refer to Appendix D.1 for more details of our evaluation tasks.
Benchmark algorithms. In our experiment, we benchmark all tree-search algorithms introduced in section 3.2.2, including MCTS-$\alpha$, MCTS-Rollout, MCTS, BFS, and DFS. We mainly compare them with direct decoding methods: CoT greedy decoding (denoted as CoT), and CoT with self-consistency (denoted as CoT-SC). Considering the search space gap between direct decoding and tree decoding (especially the sentence-level action node discussed in the section 3.1), we also include another baseline - CoT-SC-Tree which conducts CoT-SC over the tree’s sentence nodes.
@@
Model and training details. For the rollout policy used in tree-search, we use LLaMA2-7B (Touvron et al., 2023b) on three reasoning tasks, and GPT-2-small (125M) on the RLHF alignment task. All LLM will be first supervise-finetuned on the tasks’ training set, enabling their zero-shot CoT ability. Thus, CoT in our experiments refers to zero-shot greedy decoding. For value and ORM training, the data are generated by randomly sampling the SFT policy’s rollouts on the training set.
4.2 Results and Discussion
Question 1: How does TS-LLM perform in different generation tasks regarding Path@1? Does Path@1 provide a reasonable evaluation metric? (Sec. 3.2.2 and Sec. 3.4)
[…]
Method Performance(%) / # Tokens Reward / # Forward
GSM8k Game24 PrOntoQA RLHF(token-level)
CoT-greedy 41.4 $\pm$ 0.0 98 12.7 $\pm$ 0.0 76 48.8 $\pm$ 0.0 77 0.387 $\pm$ 0.0 57.8
MCTS-$\alpha$ 51.9 $\pm$ 0.6 561 63.3 $\pm$ 1.9 412 99.4 $\pm$ 0.2 190 2.221 $\pm$ 0.0 186
MCTS-Rollout 47.8 $\pm$ 0.8 3.4k 71.3 $\pm$ 2.5 670 96.9 $\pm$ 0.6 210 1.925 $\pm$ 0.0 809
MCTS 52.2 $\pm$ 1.7 486 64.0 $\pm$ 3.2 371 94.2 $\pm$ 1.6 125 -1.295 $\pm$ 0.0 61.8
BFS 52.5 $\pm$ 1.3 485 64.8 $\pm$ 2.9 369 94.4 $\pm$ 0.3 126 -1.295 $\pm$ 0.0 61.8
DFS 51.8 $\pm$ 0.6 486 66.3 $\pm$ 1.9 369 93.3 $\pm$ 0.8 126 -1.295 $\pm$ 0.0 61.8
@@ 4.2 Results and Discussion
Table 3: Performance on Game24 with different node size.
Table 4: Results of TS-LLM variants on Game24 by prompting LLaMA2-7B as the value function and the final step ORM.
Method Performance(%) / # tokens
expand by 6 expand by 20 expand by 50
MCTS-$\alpha$ 44.2 252 63.3 413 74.6 569
Rollout 44.2 346 71.3 672 79.8 838
BFS 45.9 202 64.8 373 74.9 527
CoT-SC-Tree@10 34.0 508 48.3 651 49.2 706
CoT-SC@10 - - - - 50.8 0.8k
Method Performance(%)
CoT-greedy 12.71 $\pm$ 0.0
BFS@1 9.2 $\pm$ 3.3
MCTS-$\alpha$@1 8.08 $\pm$ 0.3
Table 4: Results of TS-LLM variants on Game24 by prompting LLaMA2-7B as the value function and the final step ORM.
In Table 2, we first provide the comparison between TS-LLM variants with the CoT baseline regarding the Path@1 performance. Path@1 means each algorithm only generates one answer. TS-LLM variants generally outperform the baseline, underscoring the advantages of tree-search algorithms due to their capability to efficiently prune the search tree. However, notable exceptions are also seen in the RLHF task where MCTS, BFS, and DFS underperform. From our results, we want to highlight two things. Firstly, MCTS, BFS, and DFS generally perform better in shallow search problems (8 for GSM8K and 4 for Game24) while MCTS-$\alpha$ and MCTS-Rollout are dominant in deep search problems (15 for PrOntoQA and 64 for RLHF). Secondly, the the backward/backpropagation operation is quite important, since the first three methods—MCTS, BFS, and DFS—do not incorporate this operation and depend greedily on the value function. Note that the MCTS Path@1 also does not include the backward operation because it only happens after finding one complete path.
@@
Question 2: How do different node constructions influence the performance? (Sec. 3.1)
As discussed in Sec. 3.1), the search space for sentence-level action nodes is limited. Thus, we investigate the possible influence introduced by different tree constructions on Game24 with different node sizes. Specifically, we set the number of maximal expanded node size as 6, 20, and 50. Table 4 lists the performance and the number of tokens generated comparing TS-LLM’s variants, CoT-SC and CoT-SC-Tree. The almost doubled performance boost from 44.2 to 79.8 indicates the importance of appropriate expansion node size, improving TS-LLM’s performance upper bound.
Question 3: Why do we need a learned value function and how to train that? (Sec. 3.2.1)
Table 5: Different value training for iterative update on GSM8k
Figure 2: Different Value training on GSM8k [Uncaptioned image]
Method Policy Value Performance(%)
MCTS-$\alpha$ $\pi_{\theta_{0}}$$\{v,\hat{r}\}_{\theta_{0}}$ 51.9 $\pm$ 0.6
MCTS-$\alpha$ $\pi_{\theta_{0}}$$\{v,\hat{r}\}_{\theta_{1}}^{\text{RL}}$ 52.0 $\pm$ 0.5
MCTS-$\alpha$ $\pi_{\theta_{0}}$$\{v,\hat{r}\}_{\theta_{1}}$ 53.2 $\pm$ 0.3
MCTS-$\alpha$ $\pi_{\theta_{1}}$$\{v,\hat{r}\}_{\theta_{0}}$ 54.1 $\pm$ 0.9
MCTS-$\alpha$ $\pi_{\theta_{1}}$$\{v,\hat{r}\}_{\theta_{1}}^{\text{RL}}$ 55.2 $\pm$ 1.2
MCTS-$\alpha$ $\pi_{\theta_{1}}$$\{v,\hat{r}\}_{\theta_{1}}$ 56.5 $\pm$ 0.6
Table 5: Different value training for iterative update on GSM8k
Figure 3: mean/max reward for RLHF alignment and the best results of 3 aggregations for the rest three tasks w.r.t. number of sequences on the 1st row and number of computations on the 2nd row.
In Table 4, we provide a motivating example by prompting LLaMA2-7b-base as the value function and reward model for TS-LLM on Game24. The performance drop of BFS and MCTS-$\alpha$ indicates the incapability of small language models as reliable evaluators, which motivates us to substitute it with a learned value function and ORM as a general solution for any task and LLM with any size.
@@
Question 4: How TS-LLM is improved by aggregating over multiple results? (Sec. 3.2.3)
@@
In Fig. 3, we demonstrate the mean/max reward for the RLHF task and the best of 3 aggregation results for the rest three tasks. We measure the performance of aggregation w.r.t path number and token consumption. From the figure, we mainly summarize two conclusions: Firstly, Most TS-LLM variants benefit from aggregation and can show large strengths compared with other baselines. CoT-SC only beats TS-LLM in GSM8k with the same token size, mainly because of its larger search space. TS-LLM variants are still dominant when compared with CoT-SC-Tree. Secondly, tree-search algorithms’ aggregation benefits less than CoT-SC in small-scale problems. In GSM8K and Game24, TS-LLM struggles to improve under large aggregation numbers. We believe this is because of: (1) The search space gap between CoT-SC and tree-search algorithms. Tree-search algorithms inherently explore fewer sentences, which is validated by comparing token consumption between CoT-SC-Tree@50 and CoT-SC@50. (2) Different tree searches are not independent. The latter search might be influenced by the previous one, which decreases generation diversity.
Question 5: Can TS-LLM further train LLM iteratively? (Sec. 3.3)
Table 6: Iterative update results. $\theta_{0}$ is the old parameter while $\theta_{1}$ is the new one after one iteration. We compare all combinations of policy and value on GSM8k (left) and RLHF alignment (right).
Method Policy Value Performance(%)
Greedy $\pi_{\theta_{0}}$ 41.4 $\pm$ 0.0
Greedy $\pi_{\theta_{1}}$ 47.9 $\pm$ 0.0
Greedy RFT k=50 - 47.0 $\pm$ 0.0
Greedy RFT k=100 - 47.5 $\pm$ 0.0
MCTS-$\alpha$ $\pi_{\theta_{0}}$$\{v,\hat{r}\}_{\theta_{0}}$ 51.9 $\pm$ 0.6
MCTS-$\alpha$ $\pi_{\theta_{0}}$$\{v,\hat{r}\}_{\theta_{1}}$ 53.2 $\pm$ 0.3
MCTS-$\alpha$ $\pi_{\theta_{1}}$$\{v,\hat{r}\}_{\theta_{0}}$ 54.1 $\pm$ 0.9
MCTS-$\alpha$ $\pi_{\theta_{1}}$$\{v,\hat{r}\}_{\theta_{1}}$ 56.5 $\pm$ 0.6
Method Policy Value Performance(%)
Greedy $\pi_{\theta_{0}}$ 0.39 $\pm$ 0.0
Greedy $\pi_{\theta_{1}}$ 1.87 $\pm$ 0.0
Greedy RFT N=5 - 1.16 $\pm$ 0.0
Greedy PPO - 2.53 $\pm$ 0.0
MCTS-$\alpha$ $\pi_{\theta_{0}}$$\{v,\hat{r}\}_{\theta_{0}}$ 2.221 $\pm$ 0.0
MCTS-$\alpha$ $\pi_{\theta_{0}}$$\{v,\hat{r}\}_{\theta_{1}}$ 2.482 $\pm$ 0.0
MCTS-$\alpha$ $\pi_{\theta_{1}}$$\{v,\hat{r}\}_{\theta_{0}}$ 2.532 $\pm$ 0.0
MCTS-$\alpha$ $\pi_{\theta_{1}}$$\{v,\hat{r}\}_{\theta_{1}}$ 2.670 $\pm$ 0.0

### v2

@@ 3.2.2 Tree Search Algorithms
Given a learned value function, in this section, we present five types of tree-search algorithms. We leave detailed background, preliminaries, and comparisons in Appendix C.1.
@@
Breadth-First and Depth-First Search With Value Function-Based Tree-Pruning(BFS-V/DFS-V): These two search algorithms were adopted in ToT (Yao et al., 2023). The core idea is to utilize the value function to prune the tree for efficient search, while such pruning happens in tree breadth or depth respectively. BFS-V can be regarded as a beam-search with cumulative reward as the objective.
MCTS: This approach was adopted in RAP (Hao et al., 2023), which refers to classic MCTS(Kocsis & Szepesvári, 2006). It back-propagates the value on the terminal nodes, relying on a Monte-Carlo estimate of value, and it starts searching from the initial state node.
In addition to these algorithms adopted by ToT and RAP, we consider two new variants of AlphaZero-like tree-search.
MCTS with Value Function Approximation (named as MCTS-$\alpha$): This is the MCTS variants utilized in AlphaZero (Silver et al., 2017). Starting from the initial state, we choose the node of state $s_{t}$ as the root node and do several times of search simulations consisting of select, expand and evaluate and backup, where the leaf node value evaluated by the learned value function will be backpropagated to all its ancestor nodes. After the search, we choose an action proportional to the root node’s exponentiated visit count, i.e. $a\sim\frac{N(s_{t},a)^{1/\tau}}{\sum_{b}N(s_{t},b)^{1/\tau}}$, and move to the corresponding next state. The above iteration will be repeated until finished. MCTS-$\alpha$ has two main features. Firstly, MCTS-$\alpha$ cannot trace back to its previous states once it takes an action. So it cannot restart the search from the initial state unless multiple searches are conducted which will be discussed in Section 3.2.3. Secondly, in contrast to MCTS, MCTS-$\alpha$ utilizes a value function so it can conduct the backward operation during the intermediate steps, without the need to complete the full generation to obtain a Monte-Carlo estimate.
MCTS-Rollout: Combining the features from MCTS and MCTS-$\alpha$, we propose a new variant MCTS-Rollout for tree search. Similar to MCTS, MCTS-Rollout always starts from the initial state node. It further does the search simulations analogous to MCTS-$\alpha$, and the backup process can happen in the intermediate step with value function. It repeats the operations above until the process finds $N$ complete answers or reaches the computation limit (e.g. maximum number of tokens.) MCTS-Rollout can be seen as an offline version of MCTS-$\alpha$ so they may have similar application scope. The only difference is that MCTS-Rollout can scale up the token consumption for better performance since it always reconducts the search from the beginning.
Table 1: Task setups. The node, tree max with and tree max depth are search space parameters. Refer to Appendix D.2 and E.5 for how max tree-width and tree-depth are determined.
Search Space Hyperparameters
Task Category Train/test size Node Tree Max width Tree Max depth
GSM8k Mathematical Reasoning 7.5k / 1.3k Sentence 6 8
Game24 Mathematical Planning 1.0k / 0.3k Sentence 20 4
PrOntoQA Logical Reasoning 4.5k / 0.5k Sentence 6 15
RLHF Alignment 30k / 3k Token 50 64
Chess Endgame Decision Making 0.1M/0.6k Sentence 5 50
@@ 3.2.3 Multiple Search and Search Aggregation
Inspired by Wang et al. (2022) and Uesato et al. (2022) that LLM can improve its performance on reasoning tasks by sampling multiple times and aggregating the candidates, TS-LLM also has the potential to aggregate $N$ complete answers generated by multiple tree searches or multiple generations from one search (set BFS beam size $>1$).
When conducting multiple tree searches, we usually adopt intra-tree search setting. Intra-tree search conducts multiple tree searches on the same tree, thus the state space is exactly the same. Such a method is computationally efficient as the search tree can be reused multiple times. However, the diversity of multiple generations might decrease because the former tree search might influence the latter tree searches. Also, the search space is limited in sentence-level action space because they will be fixed once expanded across multiple tree searches.
We refer to Appendix E.7 for an alternative setting called Inter-tree Search where we allow resampling in the expansion process, and without further specification, all settings in our paper are under the intra-tree search setting. Our next step is to aggregate these search results to obtain the final answer. With a learned ORM, we consider the following three different aggregation methods:
Majority-Vote Wang et al. (2022) aggregates answers using majority vote: $f^{*}=\arg\max_{f}\sum\nolimits_{\mathbf{y}^{j}}\mathbf{1}_{\text{final\_ans}(\mathbf{y}^{j})=f}$, where $\mathbf{1}$ is the indicator function.
ORM-Max. Given an outcome reward model, the aggregation can choose the answer $f$ with maximum final reward, $f^{*}=\text{final\_ans}(\arg\max_{\mathbf{y}^{j}}\hat{r}_{\phi}(\mathbf{y}^{j}|\mathbf{x}^{j}))$.
ORM-Vote. Given an outcome reward model, the aggregation can choose the answer $f$ with the sum of reward, namely $f^{*}=\arg\max_{f}\sum\nolimits_{\mathbf{y}^{j};\text{final\_ans}(\mathbf{y}^{j})=f}\hat{r}_{\phi}(\mathbf{y}^{j}|\mathbf{x}^{j})$.
@@ 3.3 Enhancing LLM Training with Tree Search
In section 3.2 we discuss how tree-search can guide LLM’s decoding process during inference time. Such guidance leads to a better decoding strategy and improves the performance of given tasks. In other words, tree-search guidance can serve as a policy improvement operator. Based on this, we propose a new training and finetuning paradigm.
Assume we have an initial LLM policy $\pi_{\theta_{\text{old}}}$ (trained by conducting supervised finetuning over the original training set) and initial LLM value and ORM: $v_{\phi_{\text{old}}}$, $\hat{r}_{\phi_{\text{old}}}$ (trained by Equ. 1 from sampling the original training questions), we can have the following iterative process:
Policy Improvement: We conduct tree-search over training set based on $\pi_{\theta_{\text{old}}}$, $v_{\phi_{\text{old}}}$, and $\hat{r}_{\phi_{\text{old}}}$ to obtain improved generations, resulting in the augmented dataset $\mathcal{D}$ and also the filtered positive examples $\mathcal{D}^{+}$.
Policy Distillation: With the tree-search-improved dataset $\mathcal{D}^{+}$, by imitating the tree-search positive trajectories, LLM policy can be further improved to $\pi_{\theta_{\text{new}}}$ with supervised loss.
$L(\theta)=\text{E}_{(\mathbf{x}^{j},\mathbf{y}^{j})\sim\mathcal{D}^{+}}\left[-\log\pi_{\theta}(\mathbf{y}^{j}|\mathbf{x}^{j})\right].$ (2)
Policy Evaluation: We train value function $v_{\phi_{\text{new}}}$ and ORM $\hat{r}_{\phi_{\text{new}}}$ over the augumented dataset $\mathcal{D}$ under loss of Equ 1.
These three processes can be conducted cyclically to iteratively refine the LLM. Such iterative process belongs to generalized policy iteration (Sutton & Barto, 2018), which is also the procedure used in AlphaZero’s training. In our case, the training process involves finetuning three networks on the tree-search augmented dataset: (1) Policy network $\pi_{\theta}$: Use cross-entropy loss with trajectories’ tokens as target (2) Value network $v_{\phi}$: Mean squared error loss with trajectories’ temporal difference (TD) or Monte-Carlo (MC) based value estimation as target, and (3) ORM $\hat{r_{\phi}}$: Mean squared error loss with trajectories’ final reward as target.
@@ 3.4 Tree Search’s Extra Computation Burdens
Tree-search algorithms will inevitably bring in additional computation burdens, especially in the node expansion phase for calculating legal child nodes and their corresponding value. Prior methodologies, such as ToT and RAP, tend to benchmark their performance against baseline algorithms using an equivalent number of generation paths (named Path@$N$). This approach overlooks the additional computational demands of the tree-search process. We also refer readers to Appendix D.3 for discussion about the computation efficiency and engineering challenges.
A more fair comparison requires monitoring the number of tokens generated for node expansion. This provides a reasonable comparison of algorithms’ performance when operating under comparable token generation conditions. We address this issue in our experiments (Sec. 4.2.2).
@@ 4 Experiments
In this section, we conduct thorough experiments to address and analyze each subsection we mention in Section 3.
@@ 4.1 Experiment Setups
Task Setups For a given MDP, the nature of the search space is primarily characterized by two dimensions: depth and width. To showcase the efficacy of tree-search algorithms across varied search spaces, we evaluate all algorithms on five tasks with different search widths and depths, including the mathematical reasoning task GSM8k (Cobbe et al., 2021), mathematical planning task Game24 (Yao et al., 2023), logical reasoning task PrOntoQA (Saparov & He, 2022), RLHF alignment task using synthetic RLHF data (Dahoas, ), and chess endgames (Abdulhai et al., 2023). The specific task statistics and search space hyperparameters are listed in Table 1. These hyperparameters, especially max search width and search depth, are determined by our exploration experiments. They can effectively present the characteristics of a task and define its search space. Refer to Appendix E.1 for more details of our evaluation tasks.
Benchmark Algorithms. We compare ToT-GPT3.5 and TS-LLM in Table 2 to verify the effectiveness of learned value function. All tree-search algorithms will be benchmarked, including MCTS-$\alpha$, MCTS-Rollout, MCTS, BFS-V, and DFS-V. Note that BFS-V, DFS-V, and MCTS are TS-LLM’s variants instead of ToT(Yao et al., 2023) or RAP(Hao et al., 2023) baselines because we adopt a learned value function rather than prompting LLM. We compare these variants with direct decoding baselines, including CoT greedy decoding, and CoT with self-consistency (Wang et al., 2022) (denoted as CoT-SC). Considering the search space gap between direct decoding and tree decoding (especially the sentence-level action node), we include the CoT-SC-Tree baseline which conducts CoT-SC over the tree’s sentence nodes.
Model and Training Details. For the rollout policy used in tree-search, we use LLaMA2-7B (Touvron et al., 2023b) on three reasoning tasks, and GPT-2-small (125M) on the RLHF task and Chess endgame. All LLMs will be first supervise-finetuned (SFT) on the training set, enabling their zero-shot CoT ability. For value and ORM training, the data are generated by sampling the SFT policy’s rollouts on the training set. Our policy LLM and value LLM are two separate models but are adapted from the same base model. Refer to Appendix D.4 for experiments on shared models.
4.2 Results and Discussions
4.2.1 Learned Value function (Sec 3.2.1)
Table 2: ToT-BFS Path@1 results with different combinations of policy and value. LLaMA-SFT and LLaMA-V refer to the trained policy and value, LLaMA and GPT-3.5 (ToT) refer to the prompt-based model for policy or value. LLaMA-V dominates compared with prompt-based value.
Task Policy Value Accuracy(%)
GSM8K GPT-3.5 GPT-3.5 (ToT) 72.7
GPT-3.5 LLaMA-V (Ours) 74.0
LLaMA-SFT LLaMA (ToT) 37.4
LLaMA-SFT GPT-3.5 (ToT) 45.8
LLaMA-SFT LLaMA-V (Ours) 52.5
Game24 GPT-3.5 GPT-3.5 (ToT) 15.5
GPT-3.5 LLaMA-V (Ours) 19.1
LLaMA-SFT LLaMA (ToT) 9.2
LLaMA-SFT GPT-3.5 (ToT) 21.0
LLaMA-SFT LLaMA-V (Ours) 64.8
We successfully show that the learned value function can be more reliable than prompt-based GPT-3.5, even though GPT-3.5 is much stronger than LLaMA2-7B. In Table 2, we conduct BFS Path@1 comparisons with different combinations of policy and value over Game 24 and GSM8K. The policy choices include few-shot GPT-3.5 and our supervised finetuned LLaMA2-7B. For the value, we utilize prompt-based GPT-3.5/LLaMA2-7B (TOT) and our learned value function LLaMA2-V. Even though the few-shot GPT-3.5 policy is an out-of-distribution policy for LLaMA2-V’s evaluation, LLaMA2-V still presents dominant performance over prompt-based GPT-3.5/LLaMA2-7B in all settings. The phenomenon of LLM’s limited self-evaluation ability by prompting aligns with other papers (Huang et al., 2023; Stechly et al., 2023). This finding largely increases the necessity of a learned value function.
@@ 4.2 Results and Discussions
Setting Method Performance(%) / # Tokens Reward / # Forward Win Rate / # Tokens
GSM8k Game24 PrOntoQA RLHF(token-level) Chess Endgame
Path@1 CoT-greedy 41.4 98 12.7 76 48.8 92 0.318 57.8 58.14 37.4
BFS-V (Ours) 52.5 485 64.8 369 94.4 126 -1.295 61.8 67.75 402
MCTS-$\alpha$ (Ours) 51.9 561 63.3 412 99.4 190 2.221 186 96.90 797
MCTS-Rollout (Ours) 47.8 3.4k 71.3 670 96.9 210 1.925 809 98.76 615
@@
4.2.2 Performance on different algorithms (Sec 3.2.2, Sec 3.4)
With a reliable learned value function, we compare the performance of different generation methods. First, in the upper part of Table 3, we present the Path@1 results of MCTS-$\alpha$ and MCTS-Rollout compared to BFS-V (BFS-/DFS-V and MCTS degenerate to greedy value tree-search in path@1 case) and CoT-Greedy. The experiment results show that AlphaZero-like search algorithms, MCTS-$\alpha$ and MCTS-Rollout significantly outperforms the baselines in tasks where long-horizon planning matters (RLHF and Chess Endgame). When searching on shallow trees, they are robust enough to maintain comparable accuracy to the baselines.
Despite the superiority of TS-LLM, we argue that the Path@1/Path@$N$ metric may not be reasonable. We also include the number of computations used in Path@1 generation (average number of tokens in sentence-level and average number of forward computation in token-level of solving one single problem). We refer readers to the second row of Fig 2 for Path@$N$ result, with token/forward number as the x-axis. TS-LLM variants consume much more computation than CoT, making the comparison unfair.
@@
Besides, we also compare the behaviors of BFS-/DFS-V and MCTS when searching for multiple paths (aggregated by the ORM model) within a comparable range of computation consumptions. Comparing these 3 variants, MCTS is almost the best w.r.t. both performance and computation cost, this indicates the importance of value back-propogation. While comparing with the Path@1 results, MCTS-$\alpha$ and MCTS-Rollout achieve comparable accuracy in shallow-search problems (GSM8k, Game24, and ProntoQA), and dominate in deep-search ones (RLHF and Chess Endgame). It verifies the necessity of Alphazero-style intermediate value back-propagation under deep-search problems.
4.2.3 Search Aggregation (Sec 3.2.3)
In Fig. 2, we demonstrate the mean/max reward for the RLHF task and the best of 3 aggregation results for GSM8K, Game24 and ProntoQA. We measure the performance of aggregation w.r.t path number and token consumption.
@@
From the figure, we mainly summarize two conclusions: First, Most TS-LLM variants benefit from aggregation and can show large strengths compared with other baselines. CoT-SC only beats TS-LLM in GSM8k with the same token size, mainly because of its larger search space. We refer the readers to Appendix D.2 for additional results on GSM8K and ProntoQA. Second, tree-search algorithms’ aggregation benefits less than CoT-SC in small-scale problems. In GSM8K and Game24, TS-LLM struggles to improve under large aggregation numbers. We believe this is because of: (1) The search space gap between CoT-SC and tree-search algorithms. Tree-search algorithms inherently explore fewer sentences, which is validated by comparing token consumption between CoT-SC-Tree@50 and CoT-SC@50. (2) Tree-search algorithms already leverage the value function and ORM, the benefits of aggregation with the ORM again become less obvious.
In all, the scalability of tree-search aggregation is an open question that is worth further exploration in future work.
4.2.4 TS-LLM for training LLM (Sec. 3.3)
Table 4: Iterative update results. $\theta_{0},\phi_{0}$ are the old parameters while $\theta_{1},\phi_{1}$ are the new ones. TS-LLM can boost performance by training LLM policy, value, or both.
Task Method Policy Value Accuracy(%)
GSM8K Greedy $\pi_{\theta_{0}}$ 41.4
$\pi_{\theta_{1}}$ 47.9
RFT-50 - 47.0
RFT-100 - 47.5
MCTS-$\alpha$ $\pi_{\theta_{0}}$$\{v,\hat{r}\}_{\phi_{0}}$ 51.9
$\pi_{\theta_{0}}$$\{v,\hat{r}\}_{\phi_{1}}$ 53.2
$\pi_{\theta_{1}}$$\{v,\hat{r}\}_{\phi_{0}}$ 54.1
$\pi_{\theta_{1}}$$\{v,\hat{r}\}_{\phi_{1}}$ 56.5
RLHF Greedy $\pi_{\theta_{0}}$ 0.39
$\pi_{\theta_{1}}$ 1.87
RFT N=5 - 1.16
PPO - 2.53
MCTS-$\alpha$ $\pi_{\theta_{0}}$$\{v,\hat{r}\}_{\phi_{0}}$ 2.22
$\pi_{\theta_{0}}$$\{v,\hat{r}\}_{\phi_{1}}$ 2.48
$\pi_{\theta_{1}}$$\{v,\hat{r}\}_{\phi_{0}}$ 2.53
$\pi_{\theta_{1}}$$\{v,\hat{r}\}_{\phi_{1}}$ 2.67
Figure 2: Aggregation results for four tasks w.r.t. number of sequences(Path@$N$) on the 1st row and the number of tokens on the 2nd row. TS-LLM benefits from aggregation but struggles to scale in small-scale problems.